OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of one half.

OpenStudy (anonymous):

A. y - 6 = one half(x + 3) B. y = one halfx - 6 C. y + 3 = one half(x - 6) D. x - 2y = 12